# Federated Unlearning Made Practical: Seamless Integration via Negated Pseudo-Gradients

### v1

@@ Abstract
The right to be forgotten is a fundamental principle of privacy-preserving regulations and extends to Machine Learning (ML) paradigms such as Federated Learning (FL). While FL enhances privacy by enabling collaborative model training without sharing private data, trained models still retain the influence of training data. Federated Unlearning (FU) methods recently proposed often rely on impractical assumptions for real-world FL deployments, such as storing client update histories or requiring access to a publicly available dataset. To address these constraints, this paper introduces a novel method that leverages negated Pseudo-gradients Updates for Federated Unlearning (PUF). Our approach only uses standard client model updates, anyway employed during regular FL rounds, and interprets them as pseudo-gradients. When a client needs to be forgotten, we apply the negated of their pseudo-gradients, appropriately scaled, to the global model. Unlike state-of-the-art mechanisms, PUF seamlessly integrates with FL workflows, incurs no additional computational and communication overhead beyond standard FL rounds, and supports concurrent unlearning requests. We extensively evaluated the proposed method on two well-known benchmark image classification datasets (CIFAR-10 and CIFAR-100) and a real-world medical imaging dataset for segmentation (ProstateMRI), using three different neural architectures: two residual networks and a vision transformer. The experimental results across various settings demonstrate that PUF achieves state-of-the-art forgetting effectiveness and recovery time, without relying on any additional assumptions, thus underscoring its practical applicability.
Index Terms:
Federated Learning, Federated Unlearning, Client Unlearning, Machine Unlearning, Privacy.
@@ I Introduction
In today’s digital landscape, privacy has become a major concern, as reflected by the emergence of robust regulatory frameworks worldwide [1]. The European Union (EU) has consistently emphasized the importance of protecting personal data, exemplified by the introduction of the General Data Protection Regulation (GDPR) in 2016 [2]. Most recently, in May 2024, the EU enacted Regulation 2024/1183 [3], establishing the European Digital Identity Framework that empowers individuals with fine-grained control over their data. One of the key rights of these regulations is the right to be forgotten, which allows individuals to request the deletion of their previously shared data. Similar rights are central to other major privacy laws worldwide, such as the California Consumer Privacy Act (CCPA) [4] where the right to delete grants California residents the on-demand removal of personal data held by businesses.
@@
Figure 1: Performance comparison of PUF against state-of-the-art mechanisms using ResNet-18 on CIFAR-100 (non-IID case) in a 10-client setup, where one client requests unlearning. The ideal FU algorithm should (1) minimize the difference with the gold standard retrained model, across forgetting metrics such as Forget Accuracy and various MIAs (in Figure expressed as $\Delta$ Forget Accuracy and $\Delta$ MIAs), (2) ensure a short recovery phase (Recovery Rounds) and (3) preserve the generalization performance (reported in Figure as $\Delta$ Test Accuracy just after unlearning, before recovery starts). A smaller polygon represents better unlearning performances. Experimental details and additional comparisons are available in Section V.
@@
Work No Historical Information No Full Participation No Stateful Clients Task Agnostic No Auxiliary Data No Computational Overhead Simultaneous Unlearning Liu et al. [15] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ FedEraser [16] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ FUKD [17] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ PGA [18] $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ NoT [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ PUF (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
An effective mechanism should achieve unlearning while maintaining the overall performance of the global model. Although some FU techniques have started to be proposed, most of them rely on hard assumptions as they usually require maintaining additional information such as historical updates [16, 15, 17] or stateful clients [18]. Moreover, these works fail to address scenarios where multiple unlearning requests occur in the same time window. This condition cannot be overlooked as clients should be allowed to remove their contributions regardless of others’ willingness.
@@
To demonstrate the broad applicability of our approach, we implemented PUF and extensively evaluated its performance on two benchmark image classification datasets (CIFAR-10, CIFAR-100) and a real-world medical imaging dataset (ProstateMRI) for segmentation. The evaluation involved two residual networks and a vision transformer. The reported results clearly show the efficiency and effectiveness of PUF in removing client contributions to the global model across a variety of settings. Figure 1 visually compares the performance of PUF against state-of-the-art mechanisms.
@@
  • We evaluated PUF across different federated settings and compared its performance to two state-of-the-art baselines for client unlearning. We show that PUF is more effective in removing the contribution of the target clients and more efficient in recovering the expected performance.
  • We open-sourced our code to support further related research in the community. The PUF code is available for the research community at: https://github.com/alessiomora/puf_unlearning.
@@ II Related Work
Table I presents a comparative analysis of PUF against existing approaches in the field. The columns indicate whether each method does not rely on a given requirement ($\checkmark$) or does depend on it ($\times$). Our comparison highlights key aspects of FU methods, with a primary focus on whether an approach requires and leverages historical updates [15, 16, 17]. While a parameter server may have the capacity to store all client contributions across training rounds and FL processes, this requirement also introduces a critical limitation: the server must maintain the ability to link past contributions with specific clients, conflicting with the ephemeral nature of FL model updates [20]. Moreover, since historical updates are pre-stored, these methods require defining a specific set of data for removal beforehand, typically a target client’s data. This constraint prevents them from handling more general sample unlearning scenarios, where the data to be forgotten is not pre-determined, such as when a client requests the removal of a fraction of its dataset. Wu et al. [17] further assume access to auxiliary unlabeled data to mitigate performance degradation due to unlearning. However, this assumption is often impractical, as the public dataset must exhibit semantic similarity to clients’ private data for effective use.
@@
Given these limitations, we focus our comparison on more realistic FU methods. Halimi et al. [18] propose Project Gradient Ascent (PGA) to erase client influence by constraining local weight updates within an L2-norm ball around a reference model. While PGA does not require historical updates or auxiliary data—making it more suitable for FL—it introduces other constraints. Clients must be stateful, storing their latest model updates, and full participation is mandatory, as unlearning relies on a reference model obtained by removing the last unlearning client’s update from the current global model. Additionally, directly applying gradient ascent on forgetting data risks gradient explosion, as the loss function often lacks an upper bound, leading to divergence and poor generalization. To counteract this, Halimi et al. employ gradient clipping and L2-norm constraints [18]. However, these techniques require careful and extensive hyperparameter tuning, such as selecting the appropriate L2-norm radius or gradient clipping threshold. PUF eliminates the risk of uncontrolled gradient growth while maintaining the standard local training procedure. This is achieved by rescaling the client’s local update (pseudo-gradient) after training, rather than modifying the training dynamics. Recently, Khalil et al. [19] proposed NoT, a novel method aiming to achieve unlearning without impractical assumptions by implementing server-side unlearning. Specifically, NoT inverts the sign of the first layer’s global model parameters upon receiving an unlearning request. While this approach does not require explicit client-side unlearning, it fails to remove a target client’s contribution selectively — NoT produces the same unlearned model regardless of which client exits the federation. Our experimental evaluation (see Figure 1 and Section V) shows that NoT requires a longer recovery period and fails to match the forgetting performance of full model retraining.
@@
In contrast to previous methods, PUF offers several key advantages. It does not rely on historical updates or auxiliary data, does not require full client participation, and operates in a stateless manner. Moreover, PUF is task-agnostic, meaning it does not need modifications based on the specific learning task. It maintains the same computational cost as standard FedAvg rounds and requires minimal tuning, as demonstrated in our experimental results. Furthermore, PUF supports sample unlearning and can handle multiple unlearning requests simultaneously. These properties make PUF a practical and effective solution for federated unlearning, inherently aligning with the design principles of FL.
@@ III-A Federated Learning
This shows that the server’s update rule in FedAvg is equivalent to applying SGD to the pseudo-gradient $-\Delta_{t}$ with $\eta_{s}=1$. This perspective reveals that FedAvg is a specific instance of the broader FedOpt framework [21], which can use various server-side optimizers.
@@ III-B Federated Unlearning
An FU method aims to remove specific learned information from the global model $w$ while preserving the good knowledge acquired on data that should not be forgotten. FU objectives are classified based on the information the algorithm is expected to forget: sample unlearning, class unlearning [22], feature unlearning [23], and client unlearning. Our work focuses on client unlearning.
@@ III-C Terminology
  • Unlearning Client: The client that requests unlearning, also referred to as the Target Client or client $u$. The objective of an unlearning algorithm is to remove the influence of client $u$ data (forgetting data) from the global model.
  • Original Model: The global model trained with the participation of both the unlearning client and all other clients.
  • Retrained Model: The global model trained without the unlearning client from the beginning of the FL process. This serves as the gold-standard baseline for evaluating the effectiveness of unlearning methods.
@@ IV-A Preliminaries
Figure 2: Overview of PUF operating modes. In PUF-Regular Round, unlearning clients provide their model updates together with other clients. In PUF-Special Round, only unlearning clients share their model updates.
@@ IV-B Federated Unlearning via Negated Pseudo-Gradients
In FedAvg [9], clients receive the weights of the current global model and send back model updates. As discussed in Section III, these updates can be interpreted as pseudo-gradients, also enabling the use of server-side optimizers [21]. We leverage this interpretation for client unlearning in PUF. Specifically, PUF flips the sign of the model update computed by the unlearning client(s), applying it in the opposite direction of the local optimization. Figure 2 provides an illustrative overview our method.
This approach emulates the effect of gradient ascent-based unlearning techniques by pushing the global model parameters away from regions where the loss function on the forgetting data is minimized. A key technical advantage of PUF over directly applying gradient ascent—i.e., maximizing the loss function instead of minimizing it on the forgetting data—is that it avoids the risk of gradient explosion. Since the loss function generally lacks an upper bound, gradient ascent for unlearning can lead to uncontrolled gradient growth, preventing convergence and severely degrading the model’s generalization ability. To address this, methods such as weight projection (e.g., L2-norm ball) and gradient clipping are often employed to constrain parameter updates. However, these techniques require careful and often extensive hyperparameter tuning, including selecting an appropriate L2-norm radius or gradient clipping threshold.
In contrast, our method eliminates the risk of uncontrolled gradient growth by preserving the standard local training procedure on clients while simply negating and rescaling the resulting update (pseudo-gradient) after local training. PUF supports two alternative operating modes:
  • Regular Round with Modified Aggregation (PUF-Regular). In this mode, unlearning is integrated into a regular training round, where both unlearning and remaining clients participate. When one or more clients opt to leave the federation and be forgotten, the round proceeds as usual, except that the aggregated model updates from unlearning clients are negated. The sign flip can be applied either locally at clients, before transmitting back the updates, or during the server-side aggregation. The updates from other clients remain unchanged. The resulting aggregation can be formalized as follows:
@@
    with $\eta_{r}$ and $\eta_{u}$ being global learning rates. In particular, $\eta_{u}$ scales the aggregated update from unlearning clients. Considering SGD as a server-side optimizer and $\eta_{s}=1$ (see Eq. 3), the resulting unlearning model is generated as:
@@
  • Special Unlearning Round with Only Unlearning Clients (PUF-Special). In this mode, a dedicated unlearning round is conducted before resuming regular training. Only the unlearning clients participate, retrieving the latest global model and performing regular training. Assuming SGD as the server-side optimizer with $\eta_{s}=1$, the aggregated update is negated and applied to the global model as follows:
@@
    where $\eta_{u}$ is the scaling factor for the aggregated update from unlearning clients. Regular FL training resumes from $w_{\bar{u}}$.
The effectiveness of PUF depends on the magnitude of the unlearning updates, which must be carefully controlled to prevent excessive model drift. As we discuss in Section V, the tuning of the scaling factor $\eta_{u}$ is crucial for balancing unlearning effectiveness while mitigating degradation in the model’s generalization performance.
PUF efficiently removes contributions from unlearning clients while remaining fully compatible with the traditional FedAvg. After the unlearning round, the global model may require a few additional rounds to recover its generalization performance, as observed in all FU mechanisms [13]. Algorithm 1 reports a unified framework for both PUF-Special and PUF-Regular. For readability, at line 14, we assume SGD with $\eta_{s}$=1.0 as the server-side optimizer.
It is worth noting that in both operating modes, clients can transmit model updates to the server, as done in traditional FedAvg, without requiring any modifications or additional overhead. PUF is designed to ensure the unlearning process is task-agnostic and seamless for federated participants, supporting multiple concurrent unlearning requests. Furthermore, as the server performs unlearning, our method also integrates seamlessly with existing security mechanisms.
Algorithm 1 PUF Algorithm. Note that $S_{t}^{+}=\emptyset$ for PUF-Special rounds, while $S_{t}^{+}\neq\emptyset$ and $S_{t}^{-}\neq\emptyset$ for PUF-Regular rounds. $S_{t}^{-}=\emptyset$ for regular training rounds.
@@
4:     $S_{t}^{+}\leftarrow$ (random set of remaining clients)
5:     $\triangleright$ $S_{t}^{+}=\emptyset$ if PUF-Special
6:     $S_{t}^{-}\leftarrow$ (set of unlearning clients)
7:     $\triangleright$ $S_{t}^{-}=\emptyset$ if no unlearning request
8:     for each client $i\in S_{t}^{+}$ simultaneously do
9:         $\Delta_{t}^{i}\leftarrow$ ClientOpt$(i,w_{t})$
10:     for each client $j\in S_{t}^{-}$ simultaneously do
11:         $\Delta_{t}^{j}\leftarrow$ ClientOpt$(j,w_{t})$
12:     $\Delta_{t}^{+}\leftarrow\frac{1}{n}\sum_{i\in S_{t}^{+}}\Delta_{t}^{i}$
13:     $\Delta_{t}^{-}\leftarrow\frac{1}{n}\sum_{j\in S_{t}^{-}}\Delta_{t}^{j}$
14:     $\Delta_{t}\leftarrow\eta_{r}\Delta_{t}^{+}-\eta_{u}\Delta_{t}^{-}$
15:     $w_{t+1}\leftarrow w_{t}+\Delta_{t}$
16:procedure ClientOpt$(k,w_{t})$
17:     $w\leftarrow w_{t}$
18:     $\mathcal{B}\leftarrow$ (split $\mathcal{D}_{k}$ into batches of size $B$)
19:     for each local epoch $e$ from 1 to $E$ do
20:         for each batch $b\in\mathcal{B}$ do
21:              $w\leftarrow w-\eta\nabla\ell(w;b)$
22:     $\Delta_{t}^{k}\leftarrow w-w_{t}$
23:     return $\Delta_{t}^{k}$ to server
V Experimental Results
V-A Datasets, Models, and Learning Setting
We conducted a comprehensive set of experiments on image classification tasks and image segmentation tasks. For image classification, we used federated versions of the CIFAR-10 and CIFAR-100 datasets [28], which consist of 60,000 32×32 color images—50,000 for training and 10,000 for testing. CIFAR-10 comprises 10 classes, while CIFAR-100 includes 100 classes. The datasets were partitioned to simulate 10 clients, ensuring no overlapping samples among them. The experiments were performed under both Identically and Independently Distributed (IID) and non-IID settings across clients. To create non-IID data, we applied label-skew partitioning based on a distribution determined by Latent Dirichlet Allocation (LDA) [29], using concentration parameters of $\alpha=0.3$ for CIFAR-10 and $\alpha=0.1$ for CIFAR-100. Figure 3 provides a visual depiction of the label distribution among clients. For image classification, we employed a standard ResNet-18 [30] and a visual transformer, namely the MiT-B0 [31]. Before performing unlearning, we train ResNet-18 from scratch for 200 FL rounds while we fine-tune MiT-B0 for 50 FL rounds, starting from a pre-trained checkpoint. If not differently indicated, clients run one local epoch ($E=1$) for standard FedAvg.
Table II outlines the settings of our experiments. For each setting, we performed a variable amount of experiments, considering one or more clients to forget. In all experiments, we consider SGD as a server-side optimizer with $\eta_{s}$=1.0, and we set $\eta_{r}$=1.0 for PUF-Regular. If not differently indicated, we set $\eta_{u}$=2.0 for PUF-Special and $\eta_{u}$=20.0 for PUF-Regular (Section V-E discusses the tuning of our methods). Specific detail about per-setting hyper-parameters are reported in the following paragraphs.

### v2

@@ Abstract
The right to be forgotten is a fundamental principle of privacy-preserving regulations and extends to Machine Learning (ML) paradigms such as Federated Learning (FL). While FL enhances privacy by enabling collaborative model training without sharing private data, trained models still retain the influence of training data. Federated Unlearning (FU) methods recently proposed often rely on impractical assumptions for real-world FL deployments, such as storing client update histories or requiring access to a publicly available dataset. To address these constraints, this paper introduces a novel method that leverages negated Pseudo-gradients Updates for Federated Unlearning (PUF). Our approach only uses standard client model updates, which are employed during regular FL rounds, and interprets them as pseudo-gradients. When a client needs to be forgotten, we apply the negation of their pseudo-gradients, appropriately scaled, to the global model. Unlike state-of-the-art mechanisms, PUF seamlessly integrates with FL workflows, incurs no additional computational and communication overhead beyond standard FL rounds, and supports concurrent unlearning requests. We extensively evaluated the proposed method on two well-known benchmark image classification datasets (CIFAR-10 and CIFAR-100) and a real-world medical imaging dataset for segmentation (ProstateMRI), using three different neural architectures: two residual networks and a vision transformer. The experimental results across various settings demonstrate that PUF achieves state-of-the-art forgetting effectiveness and recovery time, without relying on any additional assumptions.
@@ I Introduction
In today’s digital landscape, privacy has become a major concern, as reflected by the emergence of robust regulatory frameworks worldwide [1]. The European Union (EU) has consistently emphasized the importance of protecting personal data, exemplified by the introduction of the General Data Protection Regulation (GDPR) in 2016 [2]. Most recently, in May 2024, the EU enacted Regulation 2024/1183 [3], establishing the European Digital Identity Framework that empowers individuals with fine-grained control over their information. One of the key rights of these regulations is the right to be forgotten, which allows individuals to request the deletion of their previously shared data. Similar rights are central to other major privacy laws worldwide, such as the California Consumer Privacy Act (CCPA) [4] where the right to delete grants California residents the on-demand removal of personal data held by businesses.
@@
Figure 1: Performance comparison of PUF against state-of-the-art mechanisms using ResNet-18 on CIFAR-100 (non-IID case) in a 10-client setup, where one client requests unlearning. The ideal FU algorithm should (1) minimize the difference with the gold standard retrained model, across forgetting metrics such as Forget Accuracy and various MIAs (in Figure expressed as $\Delta$ Forget Accuracy and $\Delta$ MIAs), (2) minimize computational and communication cost to recover model utility. A smaller polygon represents better unlearning performances. Experimental details are available in Section V, with full results across settings reported in Table III.
@@
Work No Historical Information No Full Participation No Stateful Clients Task Agnostic No Auxiliary Data Single-Round Unlearning Simultaneous Unlearning Selective Unlearning Liu et al. [15] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ FedEraser [16] $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ FUKD [17, 18] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ PGA [19] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ MoDe [20] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ FedAU [21] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ NoT [22] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ PUF (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
An effective mechanism should achieve unlearning while maintaining the overall performance of the global model. Although some FU techniques have started to be proposed, most of them rely on hard assumptions as they usually require maintaining additional information, such as historical updates [16, 15, 17] or stateful clients [19]. Moreover, these works fail to address scenarios where multiple unlearning requests occur in the same time window. This condition cannot be overlooked, as clients should be allowed to remove their contributions regardless of others’ willingness.
@@
To demonstrate the broad applicability of our approach, we implemented PUF and extensively evaluated its performance on two benchmark image classification datasets (CIFAR-10, CIFAR-100) and a real-world medical imaging dataset (ProstateMRI) for segmentation. The evaluation involved two residual networks and a vision transformer. The reported results clearly show the efficiency and effectiveness of PUF in removing client contributions to the global model across a variety of settings. Figure 1 visually compares the performance of PUF against five state-of-the-art mechanisms. It is worth noting that PUF can also be extended to sample unlearning with minimal modification.
@@
  • We evaluate PUF across different federated settings and compare its performance to two state-of-the-art baselines for client unlearning. We show that PUF is more effective in removing the contribution of the target clients and more efficient in recovering the expected performance.
  • We open-source our code to support further related research in the community. The PUF code is available for the research community at: https://anonymous.4open.science/r/puf_unlearning/.
@@ II Related Work
Table I presents a comparative analysis of PUF against existing approaches that target the client-unlearning case. The columns indicate whether each method does not rely on a given requirement ($\checkmark$) or does depend on it ($\times$). Our comparison highlights key aspects of FU methods, with a primary focus on whether an approach requires and leverages historical updates [15, 16, 17, 18]. Storing past client model updates raises serious scalability concerns [^1] To provide a concrete example, consider a ResNet-18 model with approximately 11 million model parameters (around 44 MB per update using 32-bit precision). Retaining updates from 1000 clients over 500 training rounds and across 10 learning tasks would require more than 220 TB of storage. Such a requirement is likely infeasible in practical FL deployments, especially as models, client counts, and training durations continue to grow [10].. Beyond storage overhead, this requirement introduces a critical limitation: the server must retain the ability to link each update to a specific client, which is fundamentally at odds with the ephemeral and privacy-preserving nature of FL model updates [23]. Furthermore, because these updates must be pre-stored, the data to be removed must be specified in advance, e.g., the entire dataset of a target client. This limitation prevents such methods from supporting more flexible unlearning scenarios, such as when a client requests to remove only a subset of its data after training has already occurred.
In particular, FedEraser [16] uses a multi-round calibration mechanism, which iteratively corrects the historical updates for a set of past retained rounds. However, for each retained round, the remaining clients in the federation must participate in the calibration step; hence, it requires clients being online and willing to participate in the calibration phase for several rounds. Wu et al. [17, 18] further assume access to auxiliary unlabeled data to mitigate performance degradation due to unlearning. However, to be effective, such a proxy dataset must be semantically aligned with the data used in the federation and balanced across classes [24]. Furthermore, assuming the availability of a public dataset, and relying on it for unlearning in FL is often unrealistic in practice [25].
@@
Given these limitations, we primarily focus our comparison on more practical FU methods. Halimi et al. [19] propose Project Gradient Ascent (PGA) to erase client influence by constraining local weight updates within an L2-norm ball around a reference model. While PGA does not require historical updates or auxiliary data, it introduces other constraints. Clients must be stateful, storing their latest model updates, and full participation is mandatory, as unlearning relies on a reference model obtained by removing the last target client’s update from the current global model. Additionally, directly applying gradient ascent on forget data risks gradient explosion, as the loss function often lacks an upper bound, leading to divergence and poor generalization. To counteract this, Halimi et al. employ gradient clipping and L2-norm constraints [19]. However, gradient-ascent techniques require careful and extensive hyperparameter tuning, such as selecting the appropriate L2-norm radius or gradient clipping threshold. PUF eliminates the risk of uncontrolled gradient growth while maintaining the standard local training procedure. This is achieved by rescaling the client’s local update (pseudo-gradient) after training, rather than modifying the training dynamics. MoDe [20] introduces a multi-round unlearning protocol alternating between degradation and memory-guidance phases. This requires additional per-round coordination among clients with respect to FedAvg, as well as more per-round computational and communication requirements. In contrast, PUF does not require any change in the regular FedAvg training round. FedAU [21] trains a local auxiliary module on forget data with randomized labels, and forms the unlearned model as a linear combination of the original model and the auxiliary module. This mechanism requires clients to train an auxiliary module during the regular FedAvg rounds, hence introducing a variation over the base protocol. Recently, Khalil et al. [22] proposed NoT, a novel method aiming to achieve unlearning without impractical assumptions by implementing server-side unlearning. Specifically, NoT inverts the sign of the first layer’s global model parameters upon receiving an unlearning request. While this approach does not require explicit client-side unlearning, it fails to remove a target client’s contribution selectively. In fact, NoT produces the same unlearned model regardless of which client exits the federation. Our experimental evaluation (see Figure 1 and Section V) shows that NoT requires a longer recovery period and fails to match the forgetting performance of full model retraining.
@@
In contrast to previous methods, PUF offers several key advantages. It does not rely on historical updates or auxiliary data, does not require full client participation, and operates in a stateless manner. Moreover, PUF is task-agnostic, meaning it does not need modifications based on the specific learning task. It maintains the same computational cost as standard FedAvg rounds and requires minimal tuning, as demonstrated in our experimental results. Furthermore, PUF supports sample unlearning and can handle multiple unlearning requests simultaneously. These properties make PUF a practical and effective solution for federated unlearning, inherently aligned with the design principles of FL.
@@ III-A Federated Learning
This shows that the server’s update rule in FedAvg is equivalent to applying SGD to the pseudo-gradient $\Delta_{t}$ with $\eta_{s}=1$. This perspective reveals that FedAvg is a specific instance of the broader FedOpt framework [26], which can use various server-side optimizers. Building on this view, Reddi et al. [26] formally treat $\Delta_{t}$ as a stochastic gradient in their FedOpt framework and prove convergence guarantees for a wide range of server optimizers (e.g., SGD, Adam, Yogi) under standard smoothness and bounded-variance assumptions. This establishes that the aggregated client update has the same mathematical role as a true gradient in the server’s optimization step, despite the server never accessing raw data.
@@ III-B Federated Unlearning
An FU method aims to remove specific learned information from the global model $w$ while preserving the good knowledge acquired on data that should not be forgotten. FU objectives are classified based on the information the algorithm is expected to forget: sample unlearning, class unlearning [27], feature unlearning [28], and client unlearning. Our work mainly focuses on client unlearning.
@@ III-C Terminology
  • target client: The client that requests unlearning, also referred to as the Target Client or client $u$. The objective of an unlearning algorithm is to remove the influence of client $u$ data (i.e., forget data) from the global model.
  • Original Model: The global model trained with the participation of both the target client and all other clients.
  • Retrained Model: The global model trained without the target client from the beginning of the FL process. This serves as the gold-standard baseline for evaluating the effectiveness of unlearning methods.
@@ IV-A Preliminaries
Figure 2: Overview of PUF operating modes. In PUF-Regular Round, target clients provide their model updates together with other clients. In PUF-Special Round, only target clients share their model updates.
@@ IV-B Federated Unlearning via Negated Pseudo-Gradients
In FedAvg [9], clients receive the weights of the current global model and send back model updates. As discussed in Section III, these updates can be interpreted as pseudo-gradients, also enabling the use of server-side optimizers [26]. This interpretation is not merely conceptual: in the FedOpt framework of Reddi et al. [26], the server update in FedAvg is mathematically equivalent to applying a gradient-based optimizer to the pseudo-gradient $\Delta_{t}$, with formal convergence guarantees under standard assumptions. We leverage this interpretation for unlearning in PUF. Specifically, PUF flips the sign of the model update (pseudo-gradient) computed by the target client(s), applying it in the opposite direction of the local optimization. Figure 2 provides an illustrative overview of our method.
This approach emulates the effect of gradient ascent-based unlearning techniques by pushing the global model parameters away from regions where the loss function on the forget data is minimized. A key technical advantage of PUF over directly applying gradient ascent, which maximizes the loss function rather than minimizing it on the forget data, is that it avoids the risk of gradient explosion. Since the loss function generally lacks an upper bound, gradient-ascent techniques can lead to uncontrolled gradient growth, preventing convergence and severely degrading the model’s generalization ability. To address this, regularization methods such as weight projection (e.g., L2-norm ball) and gradient clipping are often employed to constrain parameter updates. However, these techniques require careful and often extensive hyperparameter tuning, including selecting an appropriate L2-norm radius or gradient clipping threshold.
In contrast, our method eliminates the risk of uncontrolled gradient growth by preserving the standard local training procedure on clients while negating and rescaling the resulting update (pseudo-gradient) after local training. PUF supports two alternative operating modes:
  • Regular Round with Modified Aggregation (PUF-Regular). In this mode, unlearning is integrated into a regular training round, where both unlearning and remaining clients participate. When one or more clients opt to leave the federation and be forgotten, the round proceeds as usual, except that the aggregated model updates from target clients are negated. The sign flip can be applied either locally at clients, before transmitting back the updates, or during the server-side aggregation. The updates from other clients remain unchanged. The resulting aggregation can be formalized as follows:
@@
    with $\eta_{r}$ and $\eta_{u}$ being global learning rates. In particular, $\eta_{u}$ scales the aggregated update from target clients. Considering SGD as a server-side optimizer and $\eta_{s}=1$ (see Eq. 3), the resulting unlearning model is generated as:
@@
  • Special Unlearning Round with Only target clients (PUF-Special). In this mode, a dedicated unlearning round is conducted before resuming regular training. Only the target clients participate, retrieving the latest global model, and performing regular training. Assuming SGD as the server-side optimizer with $\eta_{s}=1$, the aggregated update is negated and applied to the global model as follows:
@@
    where $\eta_{u}$ is the scaling factor for the aggregated update from target clients. Regular FL training resumes from $w_{\bar{u}}$.
Algorithm 1 PUF Algorithm. Note that $S_{t}^{+}=\emptyset$ when PUF-Special is used for the unlearning phase, while $S_{t}^{+}\neq\emptyset$ and $S_{t}^{-}\neq\emptyset$ for PUF-Regular. $S_{t}^{-}=\emptyset$ for regular training rounds.
@@
4:  if $t$ is unlearning round then
5:   $S_{t}^{+}\leftarrow$ (random set of remaining clients)
6:   $\triangleright$ $S_{t}^{+}=\emptyset$ if PUF-Special
7:   $S_{t}^{-}\leftarrow$ (set of target clients)
[…]
9:   $S_{t}^{+}\leftarrow$ (random set of clients)
10:   $S_{t}^{-}=\emptyset$
11:  for each client $i\in S_{t}^{+}$ simultaneously do
[…]
13:  for each client $j\in S_{t}^{-}$ simultaneously do
[…]
15:  $\Delta_{t}^{+}\leftarrow\frac{1}{n}\sum_{i\in S_{t}^{+}}\Delta_{t}^{i}$
16:  $\Delta_{t}^{-}\leftarrow\frac{1}{n}\sum_{j\in S_{t}^{-}}\Delta_{t}^{j}$
17:  $\Delta_{t}\leftarrow\eta_{r}\Delta_{t}^{+}-\eta_{u}\Delta_{t}^{-}$
18:  $w_{t+1}\leftarrow w_{t}+\Delta_{t}$
[…]
20:  $w\leftarrow w_{t}$
21:  $\mathcal{B}\leftarrow$ (split $\mathcal{D}_{k}$ into batches of size $B$)
22:  for each local epoch $e$ from 1 to $E$ do
23:   for each batch $b\in\mathcal{B}$ do
24:     $w\leftarrow w-\eta\nabla\ell(w;b)$
25:  $\Delta_{t}^{k}\leftarrow w-w_{t}$
26:  return $\Delta_{t}^{k}$ to server
Algorithm 1 reports a unified framework for both PUF-Special and PUF-Regular. For readability, at line 14, we assume SGD with $\eta_{s}$=1.0 as the server-side optimizer. It is important to emphasize that PUF-Regular and PUF-Special are alternative unlearning modes, and are not meant to be used jointly or sequentially. For any given unlearning request, only one of the two routines should be applied. We included both variants to support different operational contexts: PUF-Special aligns with existing works (e.g., [19]), which isolate unlearning into dedicated rounds, while PUF-Regular allows for seamless integration into standard FL workflows.
IV-C Discussion
The effectiveness of PUF depends on the magnitude of the unlearning updates, which must be carefully controlled to prevent excessive model drift. As we discuss in Section V, the tuning of the scaling factor $\eta_{u}$ is crucial for balancing unlearning effectiveness while mitigating degradation in the model’s generalization performance. PUF efficiently removes contributions from target clients while remaining fully compatible with the traditional FedAvg. After the unlearning round, the global model may require a few additional rounds to recover its generalization performance, as observed in all approximate FU mechanisms [13]. It is worth noting that in both operating modes, clients can transmit model updates to the server, as done in traditional FedAvg, without requiring any modifications or additional overhead. PUF is designed to ensure the unlearning process is task-agnostic and seamless for federated participants, supporting multiple concurrent unlearning requests.
Model Dataset Clients Unlearning Case Data Heterogeneity Pre-trained $E$ Task # target clients ResNet-18 CIFAR-10 10 Client LDA ($\alpha$=0.3) $\times$ 1 $\mathcal{C}$ Up to one ResNet-18 CIFAR-100 10 Client LDA ($\alpha$=0.1) or IID $\times$ 1 or 10 $\mathcal{C}$ Up to two ResNet-18 CIFAR-100 10 Sample LDA ($\alpha$=0.1) or IID $\times$ 1 or 10 $\mathcal{C}$ Up to one MiT-B0 CIFAR-100 10 Client LDA ($\alpha$=0.1) 1 $\mathcal{C}$ Up to two ResUNET ProstateMRI 6 Client Real-world feature skew $\times$ 1 $\mathcal{S}$ Up to one
TABLE II: Settings of reported results. $\mathcal{C}$=classification task, $\mathcal{S}$=segmentation task. EE=local epochs during FedAvg. # target clients reports if the experiments considers a single client or multiple clients to unlearn.
Sample Unlearning. Our work mainly focuses on client unlearning. However, PUF can be readily extended to sample unlearning scenarios. In this setting, during the unlearning round, the clients that request to be forgotten use only the subset of their local dataset that must be removed when computing the model update.
Limitation. As in prior FU methods such as [16, 19, 20, 21], PUF assumes that the target clients participate one final time in the unlearning phase to generate the scaled and negated pseudo-gradient update. This design reflects a deliberate trade-off: by requiring a last participation, we completely avoid storing historical model updates or client states, preserving the ephemeral nature of FL. Moreover, in cases where the client invokes its right to be forgotten for privacy reasons, our approach allows the removal of its contribution without retaining any of its past updates, thus aligning with data minimization principles.